t3 Use the Method of Substitution to evaluate the inte 5t4+39
Answer #1

I am attaching the pictures of solution. If you find my answer helpful please press thumbs up button. Thanks

let st + 3 = x - - (5 (443) to )dt = de because that com a menn] - 20 t3 dt = da - t3 dt = dx & Now when t=0 x= st+3 - x =- x=3 when t=1 x=5(1)+3 - 5+3 - x=8 so S dt ) 0 544+3 x=3